OpenStudy (anonymous):

1. 7x/8 = 21 2. Which equation matches the statement “The product of 3 and the difference of a number and 10 is 15?" 3 + (x – 10) = 15 3x – 10 = 15 3 (x – 10) = 15 3 = 15

OpenStudy (anonymous):

@abb0t @CalebBeavers can eaither of you help me?

OpenStudy (anonymous):

1. 7x=(8)(21) x=168/7 x=24

OpenStudy (abb0t):

1. multiply by 8/7(21) 2. 3(x-10) =15

OpenStudy (anonymous):

3(x-10)=15

OpenStudy (anonymous):

thx u guys